# Geometry of Quantum Logic Gates

### v1

@@ Abstract
In this work, we investigate the geometry of quantum logic gates within the holomorphic representation of quantum mechanics. We begin by embedding the physical qubit subspace into the space of holomorphic functions that are homogeneous of degree one in each Schwinger boson pair $(z_{a_{j}},z_{b_{j}})$. Within this framework, we derive explicit closed-form differential operator representations for a universal set of quantum gates—including the Pauli operators, Hadamard, CNOT, CZ, and SWAP—and demonstrate that they preserve the physical subspace exactly. Restricting to unit-magnitude variables ($|z|=1$) reveals a toroidal space $\mathbb{T}^{2N}$, on which quantum gates act as canonical transformations: Pauli operators generate Hamiltonian flows, the Hadamard gate induces a nonlinear automorphism, and entangling gates produce correlated diffeomorphisms that couple distinct toroidal factors. Beyond the torus, the full Segal–Bargmann space carries a natural Kähler geometry that governs amplitude dynamics. Entanglement is geometrically characterized via the Segre embedding into complex projective space, while topological protection arises from the $U(1)^{N}$ fiber bundle structure associated with the Jordan–Schwinger constraint.
@@ I Introduction
The representation of quantum information in continuous phase space has long provided profound insights into the structure of quantum theory, from the Wigner–Weyl correspondence to modern formulations of quantum optics and quantum computation [1, 2, 3, 4]. In [5], the authors represented both the states and the evolution of a quantum computer in phase space using the discrete Wigner function. This framework allows us to analyze quantum algorithms—such as the Fourier transform and Grover’s search—and to examine the conditions under which quantum evolution corresponds directly to classical dynamics in phase space. The Bargmann representation provides a holomorphic (complex-analytic) realization of quantum states in which wavefunctions are entire functions of a complex variable $z\in\mathbb{C}$, related to the phase-space coordinates via $z=\frac{q+ip}{\sqrt{2}}$ (in dimensionless units). In this representation, the standard phase space of classical mechanics—spanned by position $q$ and momentum pp—is identified with the complex plane, and quantum states correspond to square-integrable holomorphic functions with respect to the Gaussian measure $e^{-|z|^{2}}\,d^{2}z$. Coherent states map to simple monomials or exponentials, and operators act as differential operators, making the Segal–Bargmann space particularly well-suited for studying bosonic systems, semiclassical limits, and analytic properties of quantum dynamics. Thus, the Bargmann representation offers a powerful bridge between Hilbert space quantum mechanics and complex phase space geometry [6, 7, 8, 9, 10, 11, 12].
@@
In this work, we construct an explicit and self-contained holomorphic representation of quantum logic gates by unifying these three transformations. We show that any $N$-qubit gate can be expressed as a differential operator acting on Segal–Bargmann space functions that are homogeneous of degree one in each Schwinger boson pair $(z_{a_{j}},z_{b_{j}})$. This representation preserves the physical subspace exactly and provides a direct link between algebraic gate operations and geometric flows on phase space.
@@
Our key contributions are threefold: (i) We derive closed-form differential operator expressions for fundamental single- and multi-qubit gates—including Pauli operators, Hadamard, SWAP, CNOT, and CZ—in terms of holomorphic variables. (ii) We analyze the induced dynamics on the toroidal space $\mathbb{T}^{2N}$ obtained by restricting to unit-magnitude variables ($|z|=1$), revealing that quantum gates act as canonical transformations with distinct geometric signatures (e.g., Hamiltonian flows for Pauli gates, nonlinear automorphisms for Hadamard). (iii) We uncover deeper geometric structures beyond the torus: the full Segal–Bargmann space carries a Kähler geometry governing amplitude dynamics; entanglement is characterized via the Segre embedding; and topological protection emerges from the $U(1)^{N}$ fiber bundle structure associated with the Jordan–Schwinger constraint. The remainder of this paper is structured as follows. In Sec. II, we present the foundational mapping from fermionic systems to holomorphic representation. We derive the homogeneity constraint that characterizes physical qubit states and show how computational basis states are encoded in holomorphic variables.
Section III presents the core result: explicit differential operator representations of universal quantum logic gates in the Segal–Bargmann space. We provide closed-form expressions for single-qubit gates (Pauli operators, Hadamard) and multi-qubit gates (SWAP, CNOT, CZ), demonstrating that all preserve the physical subspace.
In Sec. IV, we restrict to unit-magnitude variables ($|z|=1$) to reveal the toroidal space $\mathbb{T}^{2N}$, where quantum gates act as canonical transformations. We analyze Pauli gates as Hamiltonian flows, Hadamard as a nonlinear automorphism, and entangling gates as diffeomorphisms that correlate distinct toroidal factors.
Section V extends the analysis beyond the torus to uncover deeper geometric structures: the Kähler geometry governing amplitude dynamics, the Segre embedding for entanglement characterization, the fiber bundle structure underlying topological protection, and the path integral formulation for semiclassical simulation.
Finally, Sec. VI summarizes our findings and discusses implications for quantum simulation, error analysis, and geometric quantum computing.
@@ II Qubits in Holomorphic Representation
Consider a system of $N$ fermionic modes with annihilation and creation operators $c_{j}$, $c_{j}^{\dagger}$ satisfying the canonical anti-commutation relations:
$\{c_{i},c_{j}^{\dagger}\}$ $=\delta_{ij},$ (1)
$\{c_{i},c_{j}\}$ $=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0.$ (2)
The Jordan–Wigner transformation maps these to spin-$\tfrac{1}{2}$ operators (Pauli matrices) on $N$ qubits:
$c_{j}$ $=\left(\bigotimes_{k=1}^{j-1}\sigma_{k}^{x}\right)\otimes\sigma_{j}^{-},$ (3)
$c_{j}^{\dagger}$ $=\left(\bigotimes_{k=1}^{j-1}\sigma_{k}^{x}\right)\otimes\sigma_{j}^{+},$ (4)
where $\sigma_{j}^{\pm}=(\sigma_{j}^{x}\pm i\sigma_{j}^{y})/2$, and $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ are Pauli matrices acting on qubit $j$. This mapping preserves the fermionic algebra through non-local string operators. Consequently, the fermionic Fock space $\mathcal{F}$ is isomorphic to the $N$-qubit Hilbert space
$\mathcal{H}_{\mathrm{qubit}}=(\mathbb{C}^{2})^{\otimes N}.$ (5)
Each qubit (spin-$\tfrac{1}{2}$) is encoded into two bosonic modes $a_{j}$, $b_{j}$ via the Schwinger boson representation:
@@
$\sigma_{j}^{z}$ $=a_{j}^{\dagger}a_{j}-b_{j}^{\dagger}b_{j}.$ (8)
Physical states for qubit $j$ are restricted to the subspace of total occupation number one:
$\hat{n}_{j}^{\mathrm{tot}}=a_{j}^{\dagger}a_{j}+b_{j}^{\dagger}b_{j}=1.$ (9)
Thus, the full $N$-qubit space $\mathcal{H}_{\mathrm{qubit}}$ is isomorphic to the physical subspace $\mathcal{H}_{\mathrm{boson}}^{\mathrm{phys}}\subset\mathcal{F}_{\mathrm{boson}}$ of the bosonic Fock space $\mathcal{F}_{\mathrm{boson}}$ for $2N$ modes, characterized by
@@
The bosonic Fock space $\mathcal{F}_{\mathrm{boson}}$ is mapped to the Segal-Bargmann space $\mathcal{H}_{\text{SB}}$, consisting of holomorphic functions $f:\mathbb{C}^{2N}\to\mathbb{C}$ that are square-integrable with respect to the Gaussian measure
@@
where $\|\mathbf{z}\|^{2}=\sum_{j=1}^{N}(|z_{a_{j}}|^{2}+|z_{b_{j}}|^{2})$.
The Bargmann map is defined by:
@@
The physical constraint $\hat{n}_{j}^{\mathrm{tot}}=1$ translates into the homogeneity condition
@@
Hence, physical states correspond to functions homogeneous of degree one in each pair $(z_{a_{j}},z_{b_{j}})$:
@@ III Holomorphic Representation of Quantum Logic Gates
Following the construction outlined above, we now express fundamental quantum logic gates—namely the Pauli gates $X$, $Y$, $Z$, the Hadamard gate $H$, and the two-qubit SWAP gate—as differential operators acting on holomorphic functions $f(\mathbf{z})\in\mathcal{B}$ that satisfy the homogeneity constraint
@@ III.1 Single-Qubit Gates
    $(\mathrm{SWAP}_{jk}f)(\dots,z_{a_{j}},z_{b_{j}},\dots,z_{a_{k}},z_{b_{k}},\dots)$ (21)
    $=f(\dots,z_{a_{k}},z_{b_{k}},\dots,z_{a_{j}},z_{b_{j}},\dots).$
@@ III.2 Multi-Qubit Gates
  • General Controlled-$U$: For any single-qubit unitary $U=u_{00}+u_{01}X+u_{10}Y+u_{11}Z$, the controlled-$U$ gate is:
@@ IV Phasor Representation of Quantum Gates
with $\phi_{a_{j}},\phi_{b_{j}}\in[0,2\pi)$. This corresponds to fixing the magnitude $|z|=1$ while retaining the phase degree of freedom, which is sufficient to represent qubit states due to the homogeneity constraint $z_{a_{j}}\partial_{z_{a_{j}}}+z_{b_{j}}\partial_{z_{b_{j}}}=1$.
Under this restriction, the physical state space becomes a torus $\mathbb{T}^{2N}$ parameterized by the phase angles $\{\phi_{a_{j}},\phi_{b_{j}}\}_{j=1}^{N}$. The basis states correspond to:
@@ IV.1 Single-Qubit Gates in Phasor Form
    $(\mathrm{SWAP}_{jk}f)(\dots,\phi_{a_{j}},\phi_{b_{j}},\dots,\phi_{a_{k}},\phi_{b_{k}},\dots)$ (39)
    $=f(\dots,\phi_{a_{k}},\phi_{b_{k}},\dots,\phi_{a_{j}},\phi_{b_{j}},\dots).$
@@ IV.2 Geometric Interpretation on Toroidal Phase Space
The restriction to unit-magnitude holomorphic variables $z_{a_{j}}=e^{i\phi_{a_{j}}}$, $z_{b_{j}}=e^{i\phi_{b_{j}}}$ maps the physical subspace of Segal-Bargmann space $\mathcal{H}_{\text{SB}}$ onto the $2N$-torus $\mathbb{T}^{2N}=(S^{1})^{2N}$, where each $S^{1}$ corresponds to a phase angle. Quantum states are then represented as functions $f(\boldsymbol{\phi})$ on this torus, and quantum gates act as canonical transformations that preserve the symplectic structure induced by the homogeneity constraint.
  • Pauli gates as Hamiltonian flows: Each Pauli operator generates a Hamiltonian flow on the toroidal phase space $\mathbb{T}^{2N}$, governed by the symplectic form induced by the homogeneity constraint. Specifically, the torus $\mathbb{T}^{2N}$ carries a natural symplectic structure $\omega=\sum_{j=1}^{N}d\phi_{a_{j}}\wedge d\phi_{b_{j}}$, under which any real-valued function $H(\boldsymbol{\phi})$ defines a Hamiltonian vector field $X_{H}$ via $\iota_{X_{H}}\omega=dH$. The Pauli operators correspond to linear Hamiltonians:
    • For $Z_{j}$, the Hamiltonian is $H_{Z_{j}}=\phi_{a_{j}}-\phi_{b_{j}}=\Delta\phi_{j}$. The resulting flow is
      $\dot{\phi}_{a_{j}}=\frac{\partial H_{Z_{j}}}{\partial\phi_{b_{j}}}=-1,\quad\dot{\phi}_{b_{j}}=-\frac{\partial H_{Z_{j}}}{\partial\phi_{a_{j}}}=1,$
      which integrates to straight-line trajectories:
      $\phi_{a_{j}}(t)=\phi_{a_{j}}(0)-t,\quad\phi_{b_{j}}(t)=\phi_{b_{j}}(0)+t.$
      Thus, $Z_{j}$ generates uniform translation along the relative phase $\Delta\phi_{j}$, while the total phase $\Sigma\phi_{j}=\phi_{a_{j}}+\phi_{b_{j}}$ remains constant—corresponding to pure phase accumulation without population transfer.

### v2

@@ Abstract
In this work, we investigate the geometry of quantum logic gates within the holomorphic representation of quantum mechanics. We begin by embedding the physical qubit subspace into the space of holomorphic functions that are homogeneous of degree one in each Schwinger boson pair $(z_{a_{j}},z_{b_{j}})$. Within this framework, we derive explicit closed-form differential operator representations for a universal set of quantum gates, including the Pauli operators, Hadamard, CNOT, CZ, and SWAP, and demonstrate that they preserve the physical subspace exactly. Restricting to unit-magnitude variables ($|z|=1$) reveals a toroidal space $\mathbb{T}^{2N}$, on which quantum gates act as canonical transformations: Pauli operators generate Hamiltonian flows, the Hadamard gate induces a nonlinear automorphism, and entangling gates produce correlated diffeomorphisms that couple distinct toroidal factors. Beyond the torus, the full Segal–Bargmann space carries a natural Kähler geometry that governs amplitude dynamics. Entanglement is geometrically characterized via the Segre embedding into complex projective space, while topological protection arises from the $U(1)^{N}$ fiber bundle structure associated with the Jordan–Schwinger constraint.
@@ I Introduction
The representation of quantum information in continuous phase space has long provided profound insights into the structure of quantum theory, from the Wigner–Weyl correspondence to modern formulations of quantum optics and quantum computation Strocchi ; Schleich ; Zachos ; review . In phase , the authors represented both the states and the evolution of a quantum computer in phase space using the discrete Wigner function. This framework allows us to analyze quantum algorithms, such as the Fourier transform and Grover’s search, and to examine the conditions under which quantum evolution corresponds directly to classical dynamics in phase space. The Bargmann representation provides a holomorphic (complex-analytic) realization of quantum states in which wavefunctions are entire functions of a complex variable $z\in\mathbb{C}$, related to the phase-space coordinates via $z=\frac{q+ip}{\sqrt{2}}$ (in dimensionless units). In this representation, the standard phase space of classical mechanics, spanned by position $q$ and momentum $p$, is identified with the complex plane, and quantum states correspond to square-integrable holomorphic functions with respect to the Gaussian measure $e^{-|z|^{2}}\,d^{2}z$. Coherent states map to simple monomials or exponentials, and operators act as differential operators, making the Segal–Bargmann space particularly well-suited for studying bosonic systems, semiclassical limits, and analytic properties of quantum dynamics. Thus, the Bargmann representation offers a powerful bridge between Hilbert space quantum mechanics and complex phase space geometry Segal ; Bargmann ; Perelomov ; Folland ; Bishop ; Hall ; almasri ; almasri1 . The space of quantum rays naturally carries a Kähler manifold structure, enabling a geometric reformulation of quantum mechanics in which states correspond to points on a symplectic manifold equipped with a compatible Riemannian metric KahlerQC . Within this framework, observables are realized as real-valued functions and Schrödinger evolution emerges as a Hamiltonian symplectic flow, closely paralleling classical mechanics. Intrinsically quantum phenomena, such as uncertainty relations and state reduction, instead originate from the Riemannian component, which lacks a classical analogue. This perspective clarifies foundational issues including second quantization, coherent-state semiclassics, and the WKB approximation, while establishing a unified setting for dynamical generalizations of quantum theory.
Complementing this general foundation, ComplexGeoQC applies the Fubini–Study geometry to low-spin Hilbert spaces, mapping physical observables and symmetries to intrinsic geometric features. This approach yields a canonical measure of entanglement, identifies Schrödinger trajectories with Killing vector fields and their associated geometric phases, and quantifies measurement uncertainty through the gradient structure of expectation-value surfaces, enabling systematic corrections to the Heisenberg principle. By further characterizing mixed states as probability distributions whose first moments reconstruct density operators, the work establishes a comprehensive geometric framework that naturally extends standard quantum formalism.
Extending these geometric and analytic insights to bosonic systems, Chabaud et al. Chabaud employ the Segal–Bargmann holomorphic representation to bridge discrete- and continuous-variable quantum information. Their central result demonstrates that Gaussian unitary evolution decomposes exactly into an integrable Calogero–Moser dynamical system governing the motion of zeros of the holomorphic wavefunction, coupled to a conformal flow of Gaussian covariance parameters. Leveraging this correspondence, they construct a systematic non-Gaussian state hierarchy, relate multimode entanglement to the factorization properties of entire functions, and derive a unified classification of subuniversal quantum computational models that generalizes both Boson Sampling and Gaussian quantum computing.
@@
In this work, we construct an explicit and self-contained holomorphic representation of quantum logic gates by unifying two fundamental mappings: (i) the Schwinger boson representation, which encodes each spin into a pair of bosonic modes subject to a total-occupation constraint; and (ii) the Segal–Bargmann transform, which maps the bosonic Fock space to a space of holomorphic functions. We show that any $N$-qubit gate can be expressed as a differential operator acting on Segal–Bargmann space functions that are homogeneous of degree one in each Schwinger boson pair $(z_{a_{j}},z_{b_{j}})$. This representation preserves the physical subspace exactly and provides a direct link between algebraic gate operations and geometric flows on phase space.
@@
Our key contributions are threefold: (i) We derive closed-form differential operator expressions for fundamental single- and multi-qubit gates, including Pauli operators, Hadamard, SWAP, CNOT, and CZ, in terms of holomorphic variables. (ii) We analyze the induced dynamics on the toroidal space $\mathbb{T}^{2N}$ obtained by restricting to unit-magnitude variables ($|z|=1$), revealing that quantum gates act as canonical transformations with distinct geometric signatures (e.g., Hamiltonian flows for Pauli gates, nonlinear automorphisms for Hadamard). (iii) We uncover deeper geometric structures beyond the torus: the full Segal–Bargmann space carries a Kähler geometry governing amplitude dynamics; entanglement is characterized via the Segre embedding; and topological protection emerges from the $U(1)^{N}$ fiber bundle structure associated with the Jordan–Schwinger constraint.
The remainder of this paper is structured as follows. In Sec. II, we present the foundational mapping from fermionic systems to holomorphic representation. We derive the homogeneity constraint that characterizes physical qubit states and show how computational basis states are encoded in holomorphic variables. Section III presents the core result: explicit differential operator representations of universal quantum logic gates in the Segal–Bargmann space. We provide closed-form expressions for single-qubit gates (Pauli operators, Hadamard) and multi-qubit gates (SWAP, CNOT, CZ), demonstrating that all preserve the physical subspace. In Sec. IV, we restrict to unit-magnitude variables ($|z|=1$) to reveal the toroidal space $\mathbb{T}^{2N}$, where quantum gates act as canonical transformations. We analyze Pauli gates as Hamiltonian flows, Hadamard as a nonlinear automorphism, and entangling gates as diffeomorphisms that correlate distinct toroidal factors. Section V extends the analysis beyond the torus to uncover deeper geometric structures: the Kähler geometry governing amplitude dynamics, the Segre embedding for entanglement characterization, the fiber bundle structure underlying topological protection, and the path integral formulation for semiclassical simulation. Finally, Sec. VI summarizes our findings and discusses implications for quantum simulation, error analysis, and geometric quantum computing.
@@ II Qubits in Holomorphic Representation
We consider a system of $N$ qubits, where each qubit $j$ is described by the Pauli operators $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ satisfying the $\mathfrak{su}(2)$ algebra
$[\sigma_{j}^{\alpha},\sigma_{k}^{\beta}]$ $=2i\,\delta_{jk}\,\epsilon_{\alpha\beta\gamma}\,\sigma_{j}^{\gamma},$ (1)
$\{\sigma_{j}^{\alpha},\sigma_{j}^{\beta}\}$ $=2\,\delta_{\alpha\beta}\,\mathbb{I},$ (2)
[…]
Each qubit is encoded into a pair of bosonic modes $(a_{j},b_{j})$ via the Schwinger boson representation Schwinger:
@@
$\sigma_{j}^{z}$ $=a_{j}^{\dagger}a_{j}-b_{j}^{\dagger}b_{j},$ (5)
where $\sigma_{j}^{\pm}=(\sigma_{j}^{x}\pm i\sigma_{j}^{y})/2$. The bosonic operators satisfy canonical commutation relations $a_{j},a_{k}^{\dagger}]=[b_{j},b_{k}^{\dagger}]=\delta_{jk}$, with all other commutators vanishing.
Physical qubit states are restricted to the subspace of total occupation number one for each mode pair:
$\hat{n}_{j}^{\mathrm{tot}}\equiv a_{j}^{\dagger}a_{j}+b_{j}^{\dagger}b_{j}=1.$ (6)
This constraint ensures that the Schwinger mapping (3)–(5) faithfully reproduces the spin-$\tfrac{1}{2}$ algebra. Consequently, the $N$-qubit Hilbert space $\mathcal{H}_{\text{qubit}}=(\mathbb{C}^{2})^{\otimes N}$ is isomorphic to the physical subspace $\mathcal{H}_{\mathrm{boson}}^{\mathrm{phys}}$ of the bosonic Fock space $\mathcal{F}_{\text{boson}}$ for $2N$ modes, defined by
@@
The bosonic Fock space $\mathcal{F}_{\text{boson}}$ is mapped to the Segal–Bargmann space $\mathcal{H}_{\text{SB}}$, consisting of holomorphic functions $f:\mathbb{C}^{2N}\to\mathbb{C}$ that are square-integrable with respect to the Gaussian measure
@@
where $\|\mathbf{z}\|^{2}=\sum_{j=1}^{N}(|z_{a_{j}}|^{2}+|z_{b_{j}}|^{2})$. The Bargmann correspondence is defined by:
@@
The physical constraint (6) translates into a homogeneity condition on the holomorphic functions:
@@
Hence, physical qubit states correspond to functions that are homogeneous of degree one in each bosonic pair $(z_{a_{j}},z_{b_{j}})$:
@@
In this representation, the logical basis states map to simple monomials:
[…]
Superpositions and entangled states correspond to linear combinations of such monomials, with the homogeneity condition (9) ensuring that the total degree in each pair $(z_{a_{j}},z_{b_{j}})$ remains exactly one.
@@ III Holomorphic Representation of Quantum Logic Gates
Following the construction outlined above, we now express fundamental quantum logic gates, namely the Pauli gates $X$, $Y$, $Z$, the Hadamard gate $H$, and the two-qubit SWAP gate, as differential operators acting on holomorphic functions $f(\mathbf{z})\in\mathcal{B}$ that satisfy the homogeneity constraint
@@ III.1 Single-Qubit Gates
    $(\mathrm{SWAP}_{jk}f)(\dots,z_{a_{j}},z_{b_{j}},\dots,z_{a_{k}},z_{b_{k}},\dots)\\ =f(\dots,z_{a_{k}},z_{b_{k}},\dots,z_{a_{j}},z_{b_{j}},\dots).$ (20)
@@ III.2 Multi-Qubit Gates
  • General Controlled-$U$: For any single-qubit unitary $U=u_{00}\mathbb{I}+u_{01}X+u_{10}Y+u_{11}Z$, the controlled-$U$ gate is:
@@ IV Phasor Representation of Quantum Gates
with $\phi_{a_{j}},\phi_{b_{j}}\in[0,2\pi)$. This corresponds to fixing the magnitude $|z|=1$ while retaining the phase degree of freedom, which is sufficient to represent qubit states due to the homogeneity constraint $z_{a_{j}}\partial_{z_{a_{j}}}+z_{b_{j}}\partial_{z_{b_{j}}}=1$. Under this restriction, the physical state space becomes a torus $\mathbb{T}^{2N}$ parameterized by the phase angles $\{\phi_{a_{j}},\phi_{b_{j}}\}_{j=1}^{N}$. The basis states correspond to:
@@ IV.1 Single-Qubit Gates in Phasor Form
    $(\mathrm{SWAP}_{jk}f)(\dots,\phi_{a_{j}},\phi_{b_{j}},\dots,\phi_{a_{k}},\phi_{b_{k}},\dots)\\ =f(\dots,\phi_{a_{k}},\phi_{b_{k}},\dots,\phi_{a_{j}},\phi_{b_{j}},\dots).$ (38)
@@ IV.2 Geometric Interpretation on Toroidal Phase Space
The restriction to unit-magnitude holomorphic variables $z_{a_{j}}=e^{i\phi_{a_{j}}}$, $z_{b_{j}}=e^{i\phi_{b_{j}}}$ maps the physical subspace of Segal-Bargmann space $\mathcal{H}_{\text{SB}}$ onto the $2N$-torus $\mathbb{T}^{2N}=(S^{1})^{2N}$, where each $S^{1}$ corresponds to a phase angle. Quantum states are then represented as functions $f(\bm{\phi})$ on this torus, and quantum gates act as canonical transformations that preserve the symplectic structure induced by the homogeneity constraint.
IV.2.1 Hamiltonian Flow Definition
To establish the geometric framework rigorously, we first define the Hamiltonian flow structure on $\mathbb{T}^{2N}$.
Definition 1 (Hamiltonian Flow on $\mathbb{T}^{2N}$).
Let $(\mathbb{T}^{2N},\omega)$ be the $2N$-dimensional torus equipped with the canonical symplectic form
$\omega=\sum_{j=1}^{N}d\phi_{a_{j}}\wedge d\phi_{b_{j}}.$ (39)
For any smooth real-valued function $H:\mathbb{T}^{2N}\to\mathbb{R}$ (the Hamiltonian), the associated Hamiltonian vector field $X_{H}$ is uniquely defined by the relation
$\iota_{X_{H}}\omega=dH,$ (40)
where $\iota_{X_{H}}$ denotes the interior product. In local coordinates $(\bm{\phi}_{a},\bm{\phi}_{b})$, this yields Hamilton’s equations:
$\dot{\phi}_{a_{j}}=\frac{\partial H}{\partial\phi_{b_{j}}},\quad\dot{\phi}_{b_{j}}=-\frac{\partial H}{\partial\phi_{a_{j}}}.$ (41)
The Hamiltonian flow $\Phi_{H}^{t}:\mathbb{T}^{2N}\to\mathbb{T}^{2N}$ is the one-parameter family of diffeomorphisms generated by integrating $X_{H}$ over time $t$:
$\frac{d}{dt}\Phi_{H}^{t}(\bm{\phi})=X_{H}\big(\Phi_{H}^{t}(\bm{\phi})\big),\quad\Phi_{H}^{0}=\text{id}.$ (42)
Proposition 1 (Symplectic Properties).
The Hamiltonian flow $\Phi_{H}^{t}$ satisfies foundations ; arnold :
  1. 1.
    Symplectomorphism: $(\Phi_{H}^{t})^{*}\omega=\omega$ (preserves symplectic structure)
  2. 2.
    Energy Conservation: $H\circ\Phi_{H}^{t}=H$ (Hamiltonian is constant along trajectories)
  3. 3.
    Liouville’s Theorem: $\text{vol}(\Phi_{H}^{t}(U))=\text{vol}(U)$ for any measurable set $U\subset\mathbb{T}^{2N}$
  4. 4.
    Group Property: $\Phi_{H}^{t+s}=\Phi_{H}^{t}\circ\Phi_{H}^{s}$
These properties ensure that quantum gates represented as Hamiltonian flows are automatically unitary, as symplectic transformations on phase space correspond to unitary transformations in the quantum Hilbert space via the metaplectic representation.
IV.2.2 Pauli Gates as Hamiltonian Flows
Each Pauli operator generates a Hamiltonian flow on the toroidal phase space $\mathbb{T}^{2N}$, governed by the symplectic form in Eq. (39). The Pauli operators correspond to specific Hamiltonian functions: